Give a closed form for the following double sum: Justify your answer.
Answer #1

Sum of n numbers: 1+2+3+ ... + n = n(n+1) / 2

sigma sigma j

= sigma n(n+1) / 2

= n*n(n+1) / 2

= O(n3)